### v1

@@ Abstract
Effective monitoring of underwater ecosystems is crucial for tracking environmental changes, guiding conservation efforts, and ensuring long-term ecosystem health. However, automating underwater ecosystem management with robotic platforms remains challenging due to the complexities of underwater imagery, which pose significant difficulties for traditional visual localization methods. We propose an integrated pipeline that combines Visual Place Recognition (VPR), feature matching, and image segmentation on video-derived images. This method enables robust identification of revisited areas, estimation of rigid transformations, and downstream analysis of ecosystem changes. Furthermore, we introduce the SQUIDLE+ VPR Benchmark—the first large-scale underwater VPR benchmark designed to leverage an extensive collection of unstructured data from multiple robotic platforms, spanning time intervals from days to years. The dataset encompasses diverse trajectories, arbitrary overlap and diverse seafloor types captured under varying environmental conditions, including differences in depth, lighting, and turbidity. Our code is available at: https://github.com/bev-gorry/underloc.
@@ I Introduction
Reliable multi-year change detection enables long term monitoring of these dynamic and fragile ecosystems, requiring highly accurate registration to capture centimetre-level variations in habitats such as coral reefs [9]. In terrestrial and aerial monitoring scenarios, change detection benefits from GPS location priors, well-defined landmarks, and many decades of robotic vision techniques mostly tailored for these scenarios. However, these methods do not directly translate to underwater surveys, which must contend with unique challenges such as limited visibility, optical distortions, turbidity, caustics, and non-linear light attenuation across depths.
@@
Finally, we introduce the SQUIDLE+ VPR Benchmark, the first large-scale benchmark for evaluating underwater VPR, leveraging publicly available data from SQUIDLE+. Through extensive experiments, we benchmark state-of-the-art VPR techniques for underwater environments, advancing precise registration for long-term ecological monitoring and improving the reliability, scalability, and effectiveness of autonomous marine surveys for scientific research and marine conservation.
@@
Figure 2: GPS trajectories (in meters) and environmental differences across underwater datasets. The visualized trajectories depict query sequences (orange) and database sequences (blue) for four underwater datasets: (a) the Eiffel Tower (2018–2020) from the Mid-Atlantic Ridge hydrothermal vent, (b) Okinawa (2016–2017–2018), capturing mesophotic coral reef environments before and after Typhoon Trami, (c) the Tasman Fracture (2018), showcasing deep-sea benthic habitats, and (d) St Helens (2011–2013), featuring images recorded during a transition movement in shallow barren zones. To analyze the unique overlap between sequences covering the same area but following different trajectories, we include a third trajectory in the Okinawa dataset. Accompanying RGB image pairs for each trajectory illustrate corresponding locations, highlighting the significant appearance variations that challenge automated VPR systems. Notably, even correctly matched locations exhibit substantial visual differences due to variations in viewpoint, lighting conditions, and actual ecosystem changes over time, further emphasizing the complexity of long-term visual place recognition in underwater environments.
@@ II Related Works
We review image registration, 3D mapping, and VPR as key components of long-term underwater ecosystem monitoring. Traditional image registration (Section II-A) relies on GPS or multi-sensor SLAM for coarse localization before alignment refinement, while our approach uses VPR to efficiently retrieve images from monocular video for precise registration. In 3D mapping, existing methods propagate semantic information within a single trajectory, whereas we extend segmentation across multi-year datasets under varying conditions (Section II-B). Underwater VPR remains underexplored, with most techniques designed for structured terrestrial environments (Section II-C). Finally, we review underwater localization benchmarking in Section II-D.
@@ II-D Underwater Localization Benchmarking
Joshi et al. [39] proposed a Visual-Inertial SLAM pipeline, evaluated on an artificial wreck off South Carolina and in Florida’s caverns and caves. Singh et al. [40] introduced a dataset for benchmarking refractive camera model estimation, featuring a time-synchronized 5-camera/IMU setup on a remotely operated vehicle (ROV) in a controlled pool environment.
@@
In this paper, we make a step towards standardised benchmarking of VPR methods in underwater scenarios, akin to efforts in terrestrial VPR [43, 44]. Specifically, we propose to leverage sequences freely available on https://squidle.org/. SQUIDLE+ hosts the largest repository of openly accessible georeferenced marine images, enabling the construction of datasets tailored to specific scientific questions and sourced from various platforms, campaigns, and globally distributed deployments. By leveraging this vast data source, our work establishes the foundation for a dedicated underwater VPR benchmark, facilitating the development and rigorous evaluation of VPR pipelines.
@@
SQUIDLE+ Okinawa 2018 2.8 01:16 0.7 37 Mesophotic coral SOTON/UTOK OPLab AUV 5.0
@@
Tasman Fracture 2018/12/04 1.2 00:57 2.3 885 Seamount habitats CSIRO O&A MRITC Towed Stereo Camera 10.0
@@
St Helens 2013 6.5 01:48 2.9 12 Shallow barren zones IMOS AUV Sirius 3.0
@@
Eiffel Tower (Mid-Atlantic Ridge) 2020 3.9 03:18 0.9 1.7k Hydrothermal vent edifice EMSO-Azores ROV 0.5
@@ III Technical Approach
This section details our approach to enable change detection in underwater scenes using video streams from uncalibrated monocular cameras.
Given two sequential image sets, $Q$ and $D$, from a video stream, we first find the top $K$ matches in $D$ for each $Q_{i}$ using global descriptors obtained via state-of-the-art VPR techniques [25, 22, 20, 30, 32, 23]. Using LightGlue [45], we rerank these matches via the inlier count, and filter out image pairs that have a large reprojection error to avoid false matches. We then extract semantic segmentation masks from the matched images, and warp the images using the homography matrix. We use the intersection-over-union (IoU) metric to approximate location similarity.
@@ III-A Visual Place Recognition – Global Retrieval Stage
We extract image descriptors using standard VPR techniques and compute the similarity matrix by evaluating the L2 norm between all image pairs: $S(Q,D)\in\mathbb{R}^{q\times d}$.
For each query image $Q_{i}$, we select the top $K$ database images $D_{k}$ with the highest similarity scores, forming the set of initial matches:
$\mathcal{M}_{K}=\{(Q_{i},D_{k})\mid D_{k}\in\operatorname{TopK}(S(Q_{i},D),K)\},$ (1)
@@ III-B Keypoint Correspondences – Local Refinement Stage
We refine matches by using LightGlue to establish keypoint correspondences $\mathcal{K}_{qd}$ in image space between SuperPoint features in $\mathcal{M}_{K}$:
$\mathcal{K}_{q}=\{\boldsymbol{k}_{q}\mid\boldsymbol{k}_{q}\in\mathbb{R}^{3}\},\quad\mathcal{K}_{d}=\{\boldsymbol{k}_{d}\mid\boldsymbol{k}_{d}\in\mathbb{R}^{3}\},$ (2)
where $\boldsymbol{k}_{q}$ and $\boldsymbol{k}_{d}$ represent keypoints expressed in homogeneous coordinates.
@@
$\mathcal{M}_{1}=\{(Q_{i},D_{1})\mid D_{1}=\operatorname{TopK}(\text{inliers}(Q_{i},D_{k}),1)\}.$ (3)
@@ III-C Homography Estimation and Outlier Filtering
From the image correspondences $\mathcal{K}_{qd}$ in the best match subset $\mathcal{M}_{1}$, we estimate the homography matrix $H\in\mathbb{R}^{3\times 3}$ and project the keypoints from one image onto their corresponding locations in the other. The reprojection error $e_{r}$ is computed as the mean root-mean-square (RMSE) distance between keypoints projected onto the opposite image:
$(\displaystyle e_{r}=\frac{1}{2\sqrt{|\mathcal{K}_{qd}|}}\Bigg{(}$$\sqrt{\sum_{(\boldsymbol{k}_{q},\boldsymbol{k}_{p})\in\mathcal{K}_{qd}}\left\|\boldsymbol{k}_{q}-H\boldsymbol{k}_{p}\right\|^{2}}$
$+\sqrt{\sum_{(\boldsymbol{k}_{q},\boldsymbol{k}_{p})\in\mathcal{K}_{qd}}\left\|H^{-1}\boldsymbol{k}_{q}-\boldsymbol{k}_{p}\right\|^{2}}\Bigg{)},$ (4)
where $|\mathcal{K}_{qd}|$ denotes the number of matched keypoints. To account for scale variations—such as when the query image is a zoomed-in version of the database image or vice versa—we average bidirectional errors by computing transformations in both directions.
Finally, we discard image pairs with a reprojection error greater than $\chi=10$ pixels, ensuring only geometrically consistent matches are retained:
$\mathcal{M}_{\chi}=\{(Q_{i},D_{1}),\ e_{r}(Q_{i},D_{1})\leq\chi\}.$ (5)
@@
To simulate a potential change detection method for visual monitoring on the registered image pairs $\mathcal{M}_{\chi}$, we automatically extract segmentation masks for each image using Segment-Anything 2 (SAM2) [46]. For simplicity in our experiments, all segmented instances within an image are merged into a single unified mask. We then use the homography matrix to warp the masks into a common image space, enabling pixel-level comparison using pixel intersection over union (IoU) as a similarity proxy.
@@ IV Experimental Setup
Figure 3: Recall@K performance for VPR methods across underwater datasets. Experimental results illustrating the probability that a correct match appears within the top $K$ retrieved candidates. We compare six VPR methods—MixVPR, CosPlace, NetVLAD, AnyLoc, MegaLoc, and CricaVPR—against a random guesser to assess whether the retrieval results, given our GPS-based ground truth, correspond to meaningful location identification. Additionally, we include the brute-force SuperPoint approach (which performs feature matching on all possible image pairs) to establish an upper performance bound. In the Okinawa dataset, the random guesser’s performance approaches that of VPR methods for $K\gtrsim 10$ due to the lawnmower trajectory pattern (see Fig. 2), which increases image density. This allows random selection to occasionally retrieve correct matches, even without visual correspondence. While this does not compromise the GPS-based ground truth, it underscores that VPR methods are most discriminative at lower $K$ values (see Section V-A). Due to its significantly higher computational cost, we do not include the SuperPoint brute-force approach in the St Helens dataset.
@@
TABLE III: Average computation time per query (ms) for each VPR method. All one-stage VPR techniques are significantly faster than the brute-force SuperPoint approach. Our hierarchical method, combining MegaLoc and SuperPoint, is ×$\times 100$ faster than brute-force SuperPoint.
@@ IV-A VPR Baselines
To assess whether the retrieval results, based on GPS-derived ground truth, correspond to meaningful place recognition, we compare the evaluated VPR methods against a random guesser, which serves as a lower performance bound. The random guesser is implemented as a Monte Carlo experiment: for each query image $Q_{i}$, we randomly select $K$ database matches, repeating this selection $n$ times per query. The Recall@K metric is then defined as the proportion of iterations where at least one correct match was retrieved among the $K$ selected database images, normalized by the total number of trials, $|Q|\cdot n$, where $|Q|$ is the number of query images and $n$ is chosen to be sufficiently large. To establish an upper performance bound, we employ a brute-force baseline that performs exhaustive local feature matching on every query-database image pair using LightGlue. In this case, the number of keypoint matches serves as the similarity measure between image pairs.
@@ IV-B SQUIDLE+ VPR Benchmark
In this work, we propose using data from SQUIDLE+ to advance underwater VPR by leveraging a vast collection of unstructured data. This dataset originates from various robotic platforms executing diverse trajectory patterns with arbitrary overlap, spanning time differences ranging from days to years. The data includes diverse seafloor types captured under varying out-of-distribution conditions of the scene (e.g., depth, lighting, turbidity).
Table I provides details of the selected sequences, while Figure 2 illustrates the robot trajectories and presents sample RGB images depicting the scene variations.
@@
To ensure consistent processing times across datasets, we downsample the RGB images to a resolution of approximately 640$\times$480 pixels, adjusted to maintain the original aspect ratio. We include AUV sequences with GPS signals to evaluate our retrieval pipeline using an image-independent ground truth for visual localization assessment [49].
@@
Okinawa [^1] https://oceanperception.com/impact/expeditions/: This dataset captures the mesophotic coral seafloor off Sesoko Island, near Okinawa, over a two-year period using the AUV Tuna-Sand [50]. The 2018 sequence was recorded days after impact from Typhoon Trami (Paeng). Visual changes from damage to the environment hinder relocalization, emphasizing the need for a robust pipeline to document damage and recovery.
@@
Figure 4: Top row: Precision-Recall curves for Best-Single-Match VPR. The SuperPoint-H method outperforms one-stage VPR approaches, achieving performance closer to the SuperPoint brute-force approach while significantly reducing computational cost (see Table III). Bottom row: The background represents the binary ground truth in black and white, with green crosses indicating true positives and red crosses indicating false positives for the best match per query using the SuperPoint-H method. Positive matches are filtered to exclude those with a reprojection error greater than 10 pixels, corresponding to a precision of 39% for the Eiffel Tower dataset, 99% for Okinawa, 22% for Tasman Fracture, and 72% for St Helens.
We include the Eiffel Tower [^4] https://www.seanoe.org/data/00810/92226/ [31] dataset alongside the SQUIDLE+ sequences to provide a curated reference specifically designed for benchmarking long-term underwater visual localization. The dataset consists of images captured during four visits to the same hydrothermal vent edifice over a five-year period. Camera poses and a common scene geometry were estimated using navigation data and SfM. In contrast to [30], where only a small subset of images was used (${\sim}1\%$), we incorporate entire sequences for a more comprehensive evaluation.

### v2

@@ Abstract
Effective monitoring of underwater ecosystems is crucial for tracking environmental changes, guiding conservation efforts, and ensuring long-term ecosystem health. However, automating underwater ecosystem management with robotic platforms remains challenging due to the complexities of underwater imagery, which pose significant difficulties for traditional visual localization methods. We propose an integrated pipeline that combines Visual Place Recognition (VPR), feature matching, and image segmentation on images extracted from video sequences. This method enables robust identification of revisited areas, estimation of rigid transformations, and downstream analysis of ecosystem changes. Furthermore, we introduce the SQUIDLE+ VPR Benchmark—the first large-scale underwater VPR benchmark designed to leverage an extensive collection of unstructured data from multiple robotic platforms, spanning time intervals from days to years. The dataset encompasses diverse trajectories with varying overlap and diverse seafloor types captured under different environmental conditions, including differences in depth, lighting, and turbidity. Our code is available at: https://github.com/bev-gorry/underloc.
@@ I Introduction
Reliable multi-year change detection enables long-term monitoring of these dynamic and fragile ecosystems, requiring highly accurate registration to capture centimeter-level variations in habitats such as coral reefs [9]. In terrestrial and aerial monitoring scenarios, change detection benefits from GPS location priors, well-defined landmarks, and many decades of robotic vision techniques mostly tailored for these scenarios. However, these methods do not directly translate to underwater surveys, which bring unique challenges such as limited visibility, optical distortions, turbidity, caustics, and non-linear light attenuation across depths.
@@
Finally, we introduce the SQUIDLE+ VPR Benchmark, the first large-scale benchmark for evaluating underwater VPR, leveraging publicly available data from SQUIDLE+. In extensive experiments, we benchmark state-of-the-art VPR techniques for underwater environments, advancing precise registration for long-term ecological monitoring and improving the reliability, scalability, and effectiveness of marine surveys for scientific research and marine conservation.
@@
Figure 2: GPS trajectories (in meters) and environmental differences across underwater datasets. The visualized trajectories depict query sequences (orange) and database sequences (blue) for four underwater datasets: (a) the Eiffel Tower (2018–2020) from the Mid-Atlantic Ridge hydrothermal vent, (b) Okinawa (2016–2017–2018), capturing mesophotic coral reef environments before and after Typhoon Trami, (c) the Tasman Fracture (2018), showcasing deep-sea benthic habitats, and (d) St Helens (2011–2013), featuring images recorded during a transition movement in shallow barren zones. To analyze the unique overlap between sequences covering the same area but following different trajectories, we include a third trajectory in the Okinawa dataset (green). Accompanying RGB image pairs for each trajectory illustrate corresponding locations, highlighting the significant appearance variations that challenge automated VPR systems. Notably, even correctly matched locations exhibit substantial visual differences due to variations in viewpoint, lighting conditions, and actual ecosystem changes over time, further emphasizing the complexity of long-term visual place recognition in underwater environments.
@@ II Related Works
We review image registration, 3D mapping, and VPR as key components of long-term underwater ecosystem monitoring. Traditional image registration (Section II-A) relies on GPS or multi-sensor SLAM for coarse localization before alignment refinement, while our approach uses VPR to efficiently retrieve images from monocular video for precise registration. In 3D mapping, existing methods propagate semantic information within a single trajectory, whereas we extend segmentation across multi-year datasets under varying conditions (Section II-B). Underwater VPR remains underexplored, with most techniques designed for structured terrestrial environments (Section II-C). Finally, we review benchmarking in underwater localization (Section II-D).
@@ II-D Underwater Localization Benchmarking
Joshi et al. [39] proposed a Visual-Inertial SLAM pipeline, evaluated on a wreck off South Carolina and in Florida’s caverns and caves. Singh et al. [40] introduced a dataset for benchmarking refractive camera model estimation, featuring a time-synchronized 5-camera/IMU setup on a remotely operated vehicle (ROV) in a controlled pool environment.
@@
In this paper, we make a step towards standardized benchmarking of VPR methods in underwater scenarios, akin to efforts in terrestrial VPR [43, 44]. Specifically, we propose to leverage sequences freely available on https://squidle.org/. SQUIDLE+ hosts the largest repository of openly accessible georeferenced marine images, enabling the construction of datasets tailored to specific scientific questions and sourced from various platforms, campaigns, and globally distributed deployments. By leveraging this vast data source, our work establishes the foundation for a dedicated underwater VPR benchmark, facilitating the development and rigorous evaluation of VPR pipelines.
@@
SQUIDLE+
Okinawa 2018 2.8 01:16 0.7 37
Mesophotic
coral
SOTON/UTOK
OPLab AUV
5.0
@@
Tasman
Fracture
2018/12/04 1.2 00:57 2.3 885
Seamount
habitats
CSIRO O&A MRITC
Towed Stereo Camera
10.0
@@
St Helens
2013 6.5 01:48 2.9 12
Shallow
barren zones
IMOS AUV
Sirius
3.0
@@
Eiffel Tower
(Mid-Atlantic
Ridge)
2020 3.9 03:18 0.9 1.7k
Hydrothermal
vent edifice
EMSO-Azores
ROV
0.5
@@ III Technical Approach
This section details our approach to enable change detection in underwater scenes using video streams from uncalibrated monocular cameras. Given two undistorted sequential image sets, $Q$ and $D$, from a video stream, we first find the top $K$ matches in $D$ for each $Q_{i}$ using global descriptors obtained via state-of-the-art VPR techniques [25, 22, 20, 30, 32, 23]. Using LightGlue [45], we rerank these matches via the inlier count, and filter out image pairs that have a large reprojection error to avoid false matches. We then extract semantic segmentation masks from the matched images, and warp the images using the homography matrix. We use the intersection-over-union (IoU) to approximate location similarity.
@@ III-A Visual Place Recognition – Global Retrieval Stage
We extract image descriptors using standard VPR techniques and compute the similarity matrix $S\in\mathbb{R}^{|D|\times|Q|}$ where $S_{j,i}$ represents the similarity between database image $D_{j}$ and query image $Q_{i}$ using the L2 norm. For each query image $Q_{i}$, we select the top $K$ database images with the highest similarity scores, forming the set of initial matches:
$\mathcal{M}_{K,i}=\{(Q_{i},D_{k})\mid D_{k}\in\operatorname{TopK}(\{D_{j}:S_{j,i}\}_{j=1}^{|D|},K)\},$ (1)
@@ III-B Keypoint Correspondences – Local Refinement Stage
Using default parameters, we refine matches by using LightGlue to establish keypoint correspondences between SuperPoint features for each candidate pair in $\mathcal{M}_{K,i}$. For a query-database pair $(Q_{i},D_{k})$, let:
$\mathcal{C}_{i,k}=\{(\boldsymbol{p}_{q},\boldsymbol{p}_{d})\mid\boldsymbol{p}_{q}\in\mathbb{R}^{2},\boldsymbol{p}_{d}\in\mathbb{R}^{2}\}$ (2)
represent the set of keypoint correspondences between query image $Q_{i}$ and database image $D_{k}$, where $\boldsymbol{p}_{q}$ and $\boldsymbol{p}_{d}$ are corresponding keypoints in image coordinates. We refer to a keypoint correspondence retained by LightGlue under its default matching criteria as an inlier.
@@
$\mathcal{M}_{1,i}=\{(Q_{i},D^{*})\mid D^{*}=\operatorname*{arg\,max}_{D_{k}\in\mathcal{M}_{K,i}}|\mathcal{C}_{i,k}|\}$ (3)
@@ III-C Homography Estimation and Outlier Filtering
From the image correspondences $\mathcal{C}_{i,k^{*}}$ in the best match $(Q_{i},D_{k^{*}})\in\mathcal{M}_{1,i}$, we estimate the homography matrix $H\in\mathbb{R}^{3\times 3}$ and compute the bidirectional reprojection error. The reprojection error $e_{r}$ is computed as the average of the root-mean-square error (RMSE) in both transformation directions:
$e_{r}=\frac{1}{2}\Bigg($ $\sqrt{\frac{1}{|\,\mathcal{C}_{i,k^{*}}|}\sum_{(\boldsymbol{p}_{q},\boldsymbol{p}_{d})\in\mathcal{C}_{i,k^{*}}}\left\|\boldsymbol{p}_{q}-H\boldsymbol{p}_{d}\right\|^{2}}$
$+\sqrt{\frac{1}{|\,\mathcal{C}_{i,k^{*}}|}\sum_{(\boldsymbol{p}_{q},\boldsymbol{p}_{d})\in\mathcal{C}_{i,k^{*}}}\left\|H^{-1}\boldsymbol{p}_{q}-\boldsymbol{p}_{d}\right\|^{2}}\Bigg),$ (4)
where $|\,\mathcal{C}_{i,k^{*}}|$ denotes the number of matched keypoints. To account for scale variations, such as when the query image is a zoomed-in version of the database image or vice versa, we compute bidirectional reprojection errors and average them to obtain a symmetric similarity measure.
Finally, we discard image pairs with reprojection error exceeding $\chi=10$ pixels to retain only geometrically consistent matches:
$\mathcal{M}_{\chi}=\{(Q_{i},D_{k^{*}})\mid e_{r}(Q_{i},D_{k^{*}})\leq\chi\}.$ (5)
This threshold balances tolerance for minor viewpoint variations with rejection of geometrically inconsistent correspondences.
@@
To simulate a potential change detection method for visual monitoring on the registered image pairs $\mathcal{M}_{\chi}$, we automatically extract segmentation masks for each image using Segment-Anything 2 (SAM2) [46]. For simplicity in our experiments, the prompt-free automatic mask generator is used and all segmented instances within an image are merged into a single unified mask. We then use the homography matrix to warp the masks into a common image space, enabling pixel-level comparison using pixel intersection over union (IoU) as a similarity proxy.
@@ IV Experimental Setup
Figure 3: Recall@K performance for VPR methods across underwater datasets. Experimental results illustrating the probability that a correct match appears within the top $K$ retrieved candidates. We compare six VPR methods—MixVPR, CosPlace, NetVLAD, AnyLoc, MegaLoc, and CricaVPR—against a random guesser to assess whether the retrieval results, given our GPS-based ground truth, correspond to meaningful location identification. Additionally, we include the brute-force SuperPoint approach (which performs feature matching on all possible image pairs) as an alternative that directly compares local features without global retrieval. In the Okinawa dataset, the random guesser’s performance approaches that of VPR methods for $K\gtrsim 10$ due to the lawnmower trajectory pattern (see Fig. 2), which increases image density. This allows random selection to occasionally retrieve correct matches, even without visual correspondence. While this does not compromise the GPS-based ground truth, it underscores that VPR methods are most discriminative at lower $K$ values (see Section V-A). Due to its significantly higher computational cost, we do not include the SuperPoint brute-force approach on the St Helens dataset.
@@
TABLE III: Average computation time per query (ms) for each VPR method. All one-stage VPR techniques are significantly faster than the brute-force SuperPoint approach. Our hierarchical method, combining MegaLoc and SuperPoint, is $100\times$ faster than brute-force SuperPoint.
@@ IV-A VPR Baselines
To assess whether the retrieval results, based on GPS-derived ground truth, correspond to meaningful place recognition, we compare the evaluated VPR methods against a random guesser, which serves as a lower performance bound. The random guesser is implemented as a Monte Carlo experiment: for each query image $Q_{i}$, we randomly select $K$ database matches, repeating this selection $n$ times per query. The Recall@K metric is then defined as the proportion of iterations where at least one correct match was retrieved among the $K$ selected database images, normalized by the total number of trials, $|Q|\cdot n$, where $|Q|$ is the number of query images and $n$ is chosen to be sufficiently large. We also compare to a brute-force baseline that performs exhaustive local feature matching on every query-database image pair using LightGlue. In this case, the number of keypoint matches serves as the similarity measure between image pairs.
@@ IV-B SQUIDLE+ VPR Benchmark
In this work, we propose using data from SQUIDLE+ to advance underwater VPR by leveraging a vast collection of unstructured data. This dataset originates from various robotic platforms executing diverse trajectory patterns with arbitrary overlap, spanning time differences ranging from days to years. The data includes diverse seafloor types captured under varying out-of-distribution scene conditions (e.g., depth, lighting, turbidity). Table I provides details of the selected sequences, while Figure 2 illustrates the robot trajectories and presents sample RGB images depicting the scene variations.
@@
[…]
@@
Okinawa [^1] https://oceanperception.com/impact/expeditions/: This dataset captures the mesophotic coral seafloor off Sesoko Island, near Okinawa, over a two-year period using the AUV Tuna-Sand [50]. The 2018 sequence was recorded days after impact from Typhoon Trami (Paeng). Visual changes from environmental damage hinder relocalization, emphasizing the need for a robust pipeline to document damage and recovery.
@@
Figure 4: Top row: Precision-Recall curves for Best-Single-Match VPR. Our hierarchical SuperPoint-H method outperforms one-stage VPR approaches, achieving performance closer to the SuperPoint brute-force approach while significantly reducing computational cost (see Table III). Bottom row: The background represents the binary ground truth in black and white, with green crosses indicating true positives and red crosses indicating false positives for the best match per query using our SuperPoint-H method. Positive matches are filtered to exclude those with a reprojection error greater than 10 pixels, corresponding to a precision of 39% for the Eiffel Tower dataset, 99% for Okinawa, 22% for Tasman Fracture, and 72% for St Helens.
We include the Eiffel Tower [^4] https://www.seanoe.org/data/00810/92226/ [31] dataset alongside the SQUIDLE+ sequences to provide a curated reference specifically designed for benchmarking long-term underwater visual localization. The dataset consists of images captured during four visits to the same hydrothermal vent edifice over a five-year period. Camera poses and a common scene geometry were estimated using navigation data and SfM. In contrast to [30], where only ${\sim}1\%$ of the images were used, we incorporate entire sequences for a more comprehensive evaluation.